In animals, the inability to produce pigmented melanin leads to albinism, a recessive disease. Two unaffected parents decided to have three children, and their first child had albinism (genotype aa). What is the probability that the second and third children will have albinism?

show me your work
Answer #1

Answer = Since albinism is a autosomal recessive disease, affected individual have genotype aa.

PARENTS -: Aa ( Father )    × Aa ( Mother )

OFFSPRING -: AA   Aa   Aa      aa ( Affected )

PROBABILIY For 2nd children = 1/4

Probability for 3rd children = 1/4

Together probability is = 1/4 × 1/4 = 1/16 that is 1/16 × 100 =

= 6.25 %

Hope you have got your answer. Give POSITIVE FEEDBACK by clicking on thumbs up so we can serve you better.
THANK YOU.